help please and thank you

5. Prove that --> 2(n+1 - 1) for all n e Zt. 6. Prove that n < 2 for all n e Z.
Answer #1

Answers S. Let PCM: È /> 2 (duti -1) When, SI-! L-His RHS 2 (dure - 1) = 2 dias – 1). = 252-2 cl : P(1) is true. assuming thaNow, 2% +1 2k+ 4-3 VK+2 k+2 2 (k+2) -3 - 2 1672 > 2 (JWTA - Swa) See, Κ »). = 4+2 > 3 = 2√k+2 253 73 25k+2 akt) = 2(VX+2 ) >6. L2h for all the For no LH.S LRH.S For Nez- L.H.S is negative number and R.H.S is positive number L.HSC RH. S For near we w